OpenStudy (elisaneedshelp):

The line y − 6 = 2(x + 4) passes through which one of the following points? A. (−4, −6) B. (4, −6) C. (−4, 6) D. (2, 6)

jigglypuff314 (jigglypuff314):

point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1) so from y − 6 = 2(x + 4) you can match up to see that x1 = -4 and y1 = 6 so a point (x1, y1) would be...

OpenStudy (elisaneedshelp):

(-4,6) !

jigglypuff314 (jigglypuff314):

Correct! :D

OpenStudy (elisaneedshelp):

omg yay